An element crystallizes both in fcc and bcc lattices. if the density of the element in the two forms is the same, the ratio of the unit cell length of fcc to that of bcc lattice is???
 
options are 
a. (2)1/3
b.  (1/2)1/3
c. (4)1/3
d.(1/4)1/3
 

Grade:12

1 Answers

Shalet Chacko
4 years ago
Option (a) from equation d=zM /Na a^3.
Taking Z1/a^3=Z2/ a^3
Here Z1 (FCC) is 4 and Z2 (BCC) IS 2 .
Now substitute this   4 / a^ 3 (fcc) = 2/ a^3 (bcc)
                   2= a ^3( fcc)/a^3(bcc)
      Therefore (2)^1/3 = a(fcc)/a(bcc)